OpenStudy (anonymous):

y varies inversely with x, and x = 7 when y = 35. Find the inverse variation equation and use it to find the value of x when y = 50.

OpenStudy (anonymous):

y=k/x 35=k/7 => k=245 so y=245/x when y =50 50=245/x => x=245/50